# Newtonian potential from scattering amplitudes in super-renormalizable gravity

### v2

@@ Abstract
Using scattering amplitudes, we compute the coupling between a general super-renormalizable gravity and massive scalar particles. This allows us to derive the $D$-dimensional Newtonian potential at both tree-level and one-loop level—the latter containing the first calculation by using newly derived three-graviton Feynman rules. In four-dimensional spacetime, we numerically demonstrate that the Newtonian potential remains finite at the origin, providing compelling evidence that the singularity-free nature of super-renormalizable gravity persists at the one-loop level.
@@ I Introduction
Super-renormalizable gravity offers a consistent and predictive framework for ultraviolet-complete QG, preserving unitarity, covariance, and perturbative control across all energy scales Biswas et al. (2012); Modesto (2012); Briscese et al. (2014); Tomboulis (2015); Calcagni and Modesto (2015); Alexander et al. (2012); Donà et al. (2015). A defining feature of this approach is weak nonlocality—a property shared by string theory Brekke et al. (1988)—which ensures super-renormalizability while effectively eliminating pathological states. While traditional perturbative studies in super-renormalizable gravity have successfully established a finite Newtonian potential at $\mathcal{O}(G)$ Biswas et al. (2012) and explored various phenomenological implications Briscese et al. (2014); Krasnikov (1987); Modesto (2012); Tomboulis (2015); Donà et al. (2015), the theory’s behavior at $\mathcal{O}(G^{2})$ has remained a critical open question. In this work, we address this outstanding challenge by leveraging modern scattering amplitude techniques to perform the first complete one-loop computation of the Newtonian potential.
The paper is organized as follows: In Section II, we introduce the super-renormalizable gravity from the perspectives of unitarity and power counting. In Section III, we give the relation between the scattering amplitudes and the matter fields in the process of gravitational radiation from a massive scalar particle, and subsequently obtained the general expression for the higher-order metric correction. Using this method, we calculate the metric correction and Newtonian potential in $D$-dimensional spacetime at the tree-level. In Section IV, by constructing the Feynman rules for the three-graviton vertex, we investigate the loop-level correction to the Newtonian potential. We demonstrate that the potential retains well-defined singularity even at the one-loop order. In Section V, where we also draw our conclusions.
@@ II.2 Power counting in Minkowski Spacetime
III The Metric correction from Tree-Level Scattering Amplitudes
@@
These two operators are independent, and their sum is equal to $\eta_{\mu\nu}$. In the inertial frame of $k_{\mu}$, these components simplify significantly, becoming diagonal and representing the time and spatial components, respectively. As depicted in Fig. 1, the amplitude in this regime encodes the metric perturbation $h_{\mu\nu}$ generated by the point-like source and its gravitational field. This amplitude, denoted $\mathcal{M}^{\mu\nu}_{\text{vertex}}$, incorporates both the energy-momentum pseudotensor and gauge-fixing contributions. A relation derived in Jakobsen (2020) links the gauge-invariant part of the scattering amplitude to the physical pseudotensor as
@@
which gives the $D$-dimensional Newtonian potential as
@@
In particular, in the case of $D=4$, the Newtonian potential can be derived as
@@
where the error function, denoted as $\text{erf}(x)$, is defined by $\text{erf}(x)=\frac{2}{\sqrt{\pi}}\int_{0}^{x}e^{-t^{2}}\,dt$. This result aligns with earlier work in Biswas et al. (2012); Burzillà et al. (2021) obtained with linear perturbation. The error function encodes the nonlocal smearing of the gravitational interaction: at large distances $\lim_{x\to\infty}\text{erf}(x)=1$, the potential reduces to the standard Newtonian form $\frac{G_{N}m}{r}$, recovering GR in the infrared. In the short-range regime, however, the potential remains perfectly regular: $\Phi^{(G_{N})}(r\to 0)=-\frac{G_{N}mM_{*}}{\sqrt{\pi}}$. The weak nonlocality regulates the ultraviolet behavior of gravity, defusing the classical singularity at tree-level.
IV The Metric correction from One-Loop Scattering Amplitudes
@@
Furthermore, focusing on the tensor part of the integral involving the three-graviton vertex from the GR contribution $\left(V^{\alpha\beta\gamma\delta\mu\nu}_{h^{3}\text{GR}}\right)$ and two factors of $f_{\alpha\beta}$, we express the amplitude from the GR part as
@@
with $(G_{N})^{2}$ order Newtonian potential
@@
Based on the above equation, we find that the Newtonian potential at order $\mathcal{O}(G_{N}^{2})$, given in Eq. (35), is structurally governed by the spacetime dimension $D$ and four loop integrals $N_{D-1},N^{(i)}_{D-1}$ (with $i=1,2,3$), whose explicit forms are provided in Appendix B. These integrals do not admit a simple analytic expression, motivating a numerical treatment in the physically relevant case $D=4$. As displayed in Fig. 3, the computed potential is plotted against the dimensionless radial coordinate, the one-loop correction grows smoothly with distance, matching expected long-range behavior. Most notably, the potential remains completely regular as $r\to 0$, offering direct numerical evidence that the mechanism of singularity resolution, enforced by weak nonlocality, persists at the quantum level.
@@ V Summary and conclusion
In this work, we investigate the modifications to the Newtonian potential in $D$-dimensional spacetime within the framework of super-renormalizable gravity, focusing on the first- and second-order corrections to the gravitational constant $G_{N}$ through the Feynman scattering amplitudes approach. We first introduce the super-renormalizable gravity, analyzing it from two key perspectives: (1) from the viewpoint of the propagator and unitarity, where we observe that unitarity is preserved at the tree-level. and (2) from the power-counting perspective, where we identify that the primary source of divergence arises from the one-loop diagrams, thereby further confirming the super-renormalizable.
Subsequently, we examine the Newtonian potential through the Feynman scattering amplitudes, focusing on both tree-level and loop-level contributions. At tree level, the potential in $D=4$ is nonsingular and governed by an error function, thereby regularizing the classical divergence. We further performed the first one-loop calculation of the $\mathcal{O}(G^{2})$ correction, which required deriving new Feynman rules for the three-graviton vertex in the nonlocal framework. Although the one-loop result involves intricate integrals that preclude simple analytic forms, a numerical evaluation in four dimensions confirms unambiguously that the potential remains finite at the origin. This provides the first conclusive evidence that singularity resolution in super-renormalizable gravity holds at the one-loop level. Whether this remarkable property persists at higher loop orders remains an important question for future investigation.

### v3

@@ Abstract
Based on the classical limit of relativistic scattering amplitudes, we compute the coupling between a general super-renormalizable gravity and massive scalar particles. This allows us to derive the $D$-dimensional metric corrections at both tree-level and one-loop level—the latter containing the first calculation by using newly derived three-graviton Feynman rules. By introducing Newtonian potential function $\Phi\equiv\frac{h_{00}}{2}$, we reproduce at tree level, $\mathcal{O}(G)$, the well-known result proportional to the error function in four-dimensional spacetime. Furthermore, we obtain the loop-level contributions to the potential function at order $\mathcal{O}(G^{2})$ and perform a numerical analysis of its behavior at large distances. Our results indicate that, at the one-loop level, the magnitude of the potential increases gradually with distance in the asymptotic regime.
@@ I Introduction
In addition, super-renormalizable gravity offers a consistent and predictive framework for ultraviolet-complete QG, preserving unitarity, covariance, and perturbative control across all energy scales Biswas et al. (2012); Modesto (2012); Briscese et al. (2014); Tomboulis (2015); Calcagni and Modesto (2015); Alexander et al. (2012); Donà et al. (2015). A defining feature of this approach is weak nonlocality—a property shared by string theory Brekke et al. (1988)—which ensures super-renormalizability while effectively eliminating pathological states. While traditional perturbative studies in super-renormalizable gravity have successfully established a finite Newtonian potential at $\mathcal{O}(G)$ Biswas et al. (2012) and explored various phenomenological implications Briscese et al. (2014); Krasnikov (1987); Modesto (2012); Tomboulis (2015); Donà et al. (2015), the structure of loop-induced metric corrections within this class of theories has remained far less explored. In this work, we address this gap by employing modern scattering-amplitude techniques to investigate the classical limit of the relativistic amplitude at one-loop order, thereby providing a systematic analysis of the leading long-range $\mathcal{O}(G^{2})$ contributions to the effective static gravitational interaction.
The paper is organized as follows: In Section II, we introduce the super-renormalizable gravity from the perspectives of unitarity and power counting. In Section III, we give the relation between the scattering amplitudes and the matter fields in the process of gravitational radiation from a massive scalar particle, and subsequently obtained the general expression for the higher-order metric correction. Using this method, we calculate the metric correction and Newtonian potential in $D$-dimensional spacetime at the tree-level. In Section IV, by constructing the Feynman rules for the three-graviton vertex, we investigate the loop-level correction to the Newtonian potential. Furthermore, we numerically analyze the behavior of the potential function as a function of distance in the large-distance regime. In Section V, where we also draw our conclusions.
@@ II.2 Power counting in Minkowski Spacetime
III The Metric corrections from Tree-Level Scattering Amplitudes
@@
These two operators are independent, and their sum is equal to $\eta_{\mu\nu}$. In the inertial frame of $k_{\mu}$, these components simplify significantly, becoming diagonal and representing the time and spatial components, respectively. As depicted in Fig. 1, the amplitude in this regime encodes the metric perturbation $h_{\mu\nu}$ generated by the point-like source and its gravitational field. This amplitude, denoted $\mathcal{M}^{\mu\nu}_{\text{vertex}}$, incorporates both the energy-momentum pseudotensor and gauge-fixing contributions. A relation derived in Ref.Jakobsen (2020) links the gauge-invariant part of the scattering amplitude to the physical pseudotensor as
@@
where we derived the first-order correction to the metric in terms of the Newtonian constant $G_{N}$. Based on Newtonian potential $\Phi\equiv\frac{h_{00}}{2}$, the $D$ dimensional $\mathcal{O}(G_{N})$ order potential function can be written as
@@
In particular, in the case of $D=4$, the potential can be derived as
@@
where the error function, denoted as $\text{erf}(x)$, is defined by $\text{erf}(x)=\frac{2}{\sqrt{\pi}}\int_{0}^{x}e^{-t^{2}}\,dt$. This result aligns with earlier work in Biswas et al. (2012); Burzillà et al. (2021) obtained with linear perturbation. The error function encodes the nonlocal smearing of the gravitational interaction: at large distances $\lim_{x\to\infty}\text{erf}(x)=1$, the potential reduces to the standard Newtonian form $\frac{G_{N}m}{r}$, recovering GR in the infrared region.
IV The Metric corrections from One-Loop Scattering Amplitudes
@@
Furthermore, focusing on the tensor part of the integral involving the three-graviton vertex from the GR contribution $\left(V^{\alpha\beta\gamma\delta\mu\nu}_{h^{3}\text{GR}}\right)$ and factors of $f_{\alpha\beta}$, we express the amplitude from the GR part as
@@
with $(G_{N})^{2}$ order potential function
@@
Based on the above equation, we find that the Newtonian potential at order $\mathcal{O}(G_{N}^{2})$, given in Eq. (35), is structurally governed by the spacetime dimension $D$ and four loop integrals $N_{D-1},N^{(i)}_{D-1}$ (with $i=1,2,3$), whose explicit forms are provided in Appendix B. These integrals do not admit a simple analytic expression, motivating a numerical treatment in the physically relevant case $D=4$. As displayed in Fig. 3, the computed potential is plotted against the dimensionless radial coordinate, the one-loop correction grows smoothly with distance, matching expected long-range behavior. Most notably, the $\mathcal{O}(G^{2})$ contribution to the potential varies at the level of $10^{-3}$ and shows a gradual increase with distance. At this order, the potential changes sign from negative to positive, indicating a transition from an attractive to a repulsive contribution, while the tree-level $\mathcal{O}(G)$ potential remains attractive and continues to dominate the interaction.
@@ V Summary and conclusion
In this work, we investigate the modifications to the metric corrections in $D$-dimensional spacetime within the framework of super-renormalizable gravity, focusing on the first- and second-order corrections to the gravitational constant $G_{N}$ through the Feynman scattering amplitudes approach. We first introduce the super-renormalizable gravity, analyzing it from two key perspectives: (1) from the viewpoint of the propagator and unitarity, where we observe that unitarity is preserved at the tree-level. and (2) from the power-counting perspective, where we identify that the primary source of divergence arises from the one-loop diagrams, thereby further confirming the super-renormalizable.
Subsequently, we analyze the static gravitational interaction derived from relativistic scattering amplitudes, focusing on both tree-level and one-loop contributions within the classical $(\hbar\rightarrow 0)$ sector. At tree level, we derive analytic expressions for the metric corrections and the associated potential function in $D$-dimensional spacetime. We find that both quantities are governed by the upper incomplete gamma function $\Gamma\left(\frac{D-3}{2},\frac{1}{4}M_{\ast}^{2}r^{2}\right)$, where $\Gamma(a,x)$ denotes the upper incomplete gamma function of order a=(D3)/2a=(D-3)/2. In four dimensions, our results reproduce the known linearized solutions of super-renormalizable gravity $\Phi^{(G_{N})}(r)=-\frac{G_{N}m}{r}\text{erf}\left(\frac{M_{\ast}r}{2}\right)$. Since our analysis focuses on the large-distance regime, we further observe that in the limit of large radial separation $r$, the potential smoothly approaches the behavior predicted by GR.
At one-loop order, we derive the $\mathcal{O}(G^{2})$ corrections to the metric components and to the associated potential function in $D$-dimensional spacetime. The loop contributions involve nontrivial vertex structures, denoted by $V^{\alpha\beta\gamma\delta\mu\nu}_{h^{3}\text{GR}}$ and $V^{\alpha\beta\gamma\delta\mu\nu}_{h^{3}\text{Nonlocal}}$, whose explicit analytic expressions are presented in the Appendix A. To organize the calculation, we introduce three classes of integrals that encode the nonanalytic dependence on the momentum transfer $q$ responsible for the long-range classical behavior of the interaction. Owing to the complexity of these integrals, a analytic evaluation is not available, and we therefore perform a numerical analysis in four-dimensions. The numerical results indicate that the $\mathcal{O}(G^{2})$ contribution to the potential varies at the level of $10^{-3}$ and increases gradually with the radial distance $r$. Moreover, at this order the potential exhibits a change of sign from negative to positive values, corresponding to a transition from an attractive to a repulsive contribution, while the tree-level $\mathcal{O}(G)$ potential remains attractive and continues to dominate the overall interaction.
Additionally, we emphasize that the present analysis is restricted to the classical limit of relativistic scattering amplitudes, which isolates the nonanalytic contributions responsible for the long-range behavior of the effective static gravitational interaction. As a consequence, the ultraviolet structure of super-renormalizable gravity and the detailed behavior of the interaction at $r=0$ are not fully addressed in this paper. These aspects are expected to be governed by analytic, contact-term contributions in momentum space, which lie beyond the scope of the amplitude-based classical sector considered here. A systematic investigation of such analytic terms and their implications for the ultraviolet behavior and potential singularity structure of the theory will be the focus of future work.